If the sides of a square are decreased by 2cm, the area is decreased by 36 cm^2. What were the dimensions of the original square?

Thank you

Answer by scott8148(6628)

let s equal the original side length ... (s-2)^2=s^2-36 ... s^2-4s+4=s^2-36 ... -4s=-40 ... s=10